# EvidenceMap: Unleashing the Power of Small Language Models with Evidence Analysis for Biomedical Question Answering

### v1

@@ Abstract
Current LLM-based approaches improve question answering performance by leveraging the internal reasoning abilities of models or incorporating external knowledge. However, when humans address professional problems, it is essential to explicitly analyze the multifaceted relationships from multiple pieces and diverse sources of evidence to achieve better answers. In this study, we propose a novel generative question answering framework for the biomedical domain, named EvidenceMap, which explicitly learns and incorporates evidence analysis with small language models (SLMs). The framework describes an evidence map for each question and fully utilizes an SLM to derive the representation of the supportive evaluation, the logical correlation, and the summarization of the related evidence, which facilitates an analysis-augmented generation with another SLM in an autoregressive way. Extensive experiments have shown that introducing an evidence analysis learning process can significantly outperform larger models and popular LLM reasoning methods.
EvidenceMap: Unleashing the Power of Small Language Models with Evidence Analysis for Biomedical Question Answering
Chang Zong, Jian Wan, Lei Zhang thanks: Corresponding author. School of Information and Electronic Engineering Zhejiang University of Science and Technology zongchang@zju.edu.cn
@@ 1 Introduction
Answering specialized questions often requires reference to a wider range of professional information, such as academic literature and expert knowledge, compared to general domain inquiries Louis et al. (2024); Singhal et al. (2025). Generally, as shown in Figure 1, humans require the integration of evidence from multiple sources to solve professional problems Lu et al. (2022). This process involves a series of analytical steps, including evidence evaluation Watts et al. (2009), evidence correlation Chen et al. (2021), and evidence summarization Hernes et al. (2015), which ultimately leads to a well-reasoned conclusion.
Figure 1: A diagram illustrates the evidence analysis process when addressing professional questions.
To address the generative question answering task supported by multiple pieces of evidence, existing modern methods center on large language models (LLMs), employing implicit reasoning over multiple relevant pieces of knowledge or explicitly generating multiple thinking steps to mitigate model hallucinations and achieve more accurate answer generation, such as the families of chain-of-thought (CoT) Wei et al. (2022); Lightman et al. (2023); Yao et al. (2024) and retrieval augmented generation (RAG) Lewis et al. (2020); Asai et al. (2023); He et al. (2024); Wang et al. (2024), respectively. However, as illustrated in Figure 2, the reasoning frameworks mentioned above overlook the explicit definition of analytical stages and the implicit learning of analytical skills, failing to accurately simulate the human solving process, perpetuating hallucinations and error propagation problems, especially in professional domains.
Figure 2: The paradigm of our framework (right) differs from methods such as CoT and RAG by employing SLMs for learning evidence analysis expicitly, thereby reducing error propagation during reasoning.
To further alleviate the aforementioned issues, this study focuses on biomedical question answering and proposes a novel framework, named EvidenceMap. The framework enables explicit learning and utilization of evidence analysis with small language models (SLMs), thus efficiently simulating the human-solving process. Specifically, our method first acquires multiple pieces of evidence from different sources, and then constructs an evidence map, utilizing a small pre-trained language model (S-PLM) to learn latent representations for evidence evaluation, evidence correlation, and evidence summarization, which serve as the output of the evidence analysis. Finally, these representations, along with textual input, are delivered into a small generative language model (S-GLM) to autoregressively derive answers. The experimental results demonstrate that using SLMs for learning and employing evidence analysis is a promising generative question-answering approach.
@@
  • We propose EvidenceMap, a novel framework for generative biomedical question answering, which mimics the analysis of evidence from humans. The framework consists of the acquisition, evaluation, correlation, and summarization of evidence and improves the efficacy of utilizing multiple pieces of evidence.
  • We implement an efficient method for learning evidence analysis in latent space, which requires only updating a small pre-trained language model and subsequently driving another small generative model to achieve better performance in generative question answering.
  • Extensive evaluations demonstrate that our framework for learning explicit evidence analysis can significantly improve the quality of responses compared to larger language models and popular LLM-based reasoning methods.
@@ 2.1 Evidence Map Modeling
We construct an evidence map for each input question to facilitate reasoning during generation. An evidence map introduces three key types of information, including the evaluation of how the evidence supports the question ($\mathcal{R}_{eval}$), the correlation between the evidence ($\mathcal{R}_{cor}$), and the summarization of all the evidence ($E_{sum}$). Generally, an evidence map can be formulated as follows.
@@
where $Q$ is the question to be addressed, $Agg(\cdot)$ is the aggregation function to derive the summarization $E_{sum}$, $\mathcal{E}_{s}$ represents a set of evidence from the source $s$, which is denoted as follows.
@@
2.2 Learning Generative QA with SLM
@@
where $f_{plm}(\theta^{*})$ is a tiny pre-trained model (e.g. BERT-base), $f_{glm}$ is a small generative model (e.g. Llama-3.2-3B) for providing the best answer $A^{*}$. Then, our learning process is to find the optimized parameter set $\theta^{*}$, which leads to a minimum generative loss for all question-answer pairs, given the corresponding evidence map $EM_{i}$ for each question $Q_{i}$. This can be denoted as below.
@@ 3 Framework
The EvidenceMap framework is illustrated in Figure 3, in which a pre-trained model is fine-tuned to learn the analysis of evidence to facilitate a small generative model that provides better responses. The three main steps are described below.
@@ 3.1 Evidence Acquisition
For each input question, we first gather relevant original evidence from multiple sources, including the intrinsic knowledge of an LLM and the external knowledge of academic papers. Specifically, for the evidence of an LLM, we ask GPT-4o Hurst et al. (2024) to take advantage of its knowledge to provide a series of key points that support the question, described in a concise language (see Appendix A.2). In addition, we extract paper snippets from existing biomedical QA datasets such as BioASQ Krithara et al. (2023) and PubMedQA Jin et al. (2019), which have already been annotated by human experts, as evidence related to a question. Therefore, an evidence-augmented QA dataset is constructed, in which each question corresponds to one piece of LLM evidence $E_{lm}$ and multiple pieces of paper evidence from $E_{p}^{1}$ to $E_{p}^{n}$. The input data can be described as follows.
@@
3.2 Latent Evidence Analysis
The Evidence Map.
As defined in Section 2.1, each input question, along with its associated evidence, is modeled as an evidence map. In addition to the evidence pieces themselves, the evidence map also incorporates the analysis among the evidence, including the evaluation of how each piece of evidence supports the question $R_{eval}^{i}$, the correlation between any two pieces of evidence $R_{cor}^{(i,j)}$, and the summarization of all the evidence $E_{sum}$.
SLM-based Evidence Analysis.
We prompt a small pre-trained language model to perform the multifaceted evidence analysis in latent space. Specifically, for evidence summarization, we first obtain the overall semantics of each evidence node, and then use a multi-layer perceptron (MLP) to aggregate and produce a summarized representation of the evidence, which is formulated as below.
@@ 3.2 Latent Evidence Analysis
where $X$ is the concatenated feature of all evidence nodes, denoted as $[X_{lm},X_{p}^{1},\cdots,X_{p}^{n}]$, $H_{s}^{(n)}$ is the embedding of the $n$-th layer. Inspired by PromptBERT Jiang et al. (2022) and PromptEOL Jiang et al. (2024), we prompt the SLM $f_{plm}$ and get the last hidden states of the model as the evidence feature $X_{p}^{i}$, which is described as follows.
@@
Meanwhile, the evaluation of how an evidence node supports the question $Q$ is derived from the same SLM by using $\texttt{Prompt}_{e}$ as below.
@@ 3.3 Analysis-Augmented Generation
With the result of the analysis of an evidence map, the final answer to the question is provided in an autoregressive way by a generative language model $f_{glm}$. We first convert all textual evidence and the question into an embedding $H_{t}$ using a text embedder, which is the first embedding layer of $f_{glm}$ as follows.
@@ 3.4 Training Objective
The autoregressive training objective concentrates on training the parameter set of SLM $\theta$ to provide better analysis representations to predict subsequent tokens accurately by the generative language model. We calculate the probability of generating the $i$-th token of the target answer $A$ with the overall prompt and all previous tokens as follows.
@@ 4.1 Implementation Details
In the step of learning evidence analysis, we employ DistilBERT Sanh (2019), a lighter but faster model distilled from BERT Devlin et al. (2019), as the SLM to be updated. In the generation step, the 1B and 3B versions of Llama-3.2 Dubey et al. (2024) are used to provide answers. We conduct experiments on two popular biomedical datasets, BioASQ Krithara et al. (2023) and PubMedQA Jin et al. (2019). During training, we set the learning rate as 5e-4, the batch size as 4, and the maximum number of paper evidence per question as 5. For each experiment, we train the pre-trained SLM for 8 epochs in a single NVIDIA A6000 GPU and evaluate the model on test sets.
@@ 4.3 Main Results
4o-mini-RAG - 0.1549 0.5848 0.7339 0.1328 0.5501 0.5874
Inference-LLM Llama3.1-RAG 8B 0.1936 0.6077 0.7401 0.1835 0.6431 0.6440
OpenBioLLM 8B 0.2196 0.5981 0.7655 0.1936 0.5726 0.6276
Llama3.2-RAG 3B 0.2956 0.6225 0.7643 0.1942 0.6102 0.6361
@@
Tuning-SLM RAG-LoRA 3B 0.3987 0.6834 0.7440 0.2443 0.6454 0.6342
EvidenceMap 1B 0.4045 0.7075 0.7511 0.2548 0.6513 0.6375
EvidenceMap 3B 0.4529 0.7387 0.7905 0.2931 0.6783 0.6610
@@
Second, Learning evidence analysis unleashes the power of SLMs to efficiently resolve biomedical questions. Compared to larger models, including the RAG method with general LLMs (GPT-4o-mini, Llama-3.1-8B) and biomedical LLM (OpenBioLLM-8B), our approach demonstrates superior performance even with a 1B model. This indicates that evidence analysis can effectively unleash the capabilities of SLMs.
@@ 4.4 Ablation Study on Evidence Analysis
w/o Eval 0.7181 $\downarrow$ 2.79% 0.7712 $\downarrow$ 2.44%
w/o Cor 0.7072 $\downarrow$ 4.26% 0.7653 $\downarrow$ 3.19%
w/o Sum 0.7239 $\downarrow$ 2.00% 0.7856 $\downarrow$ 0.62%
@@ 4.5 Ablation Study on Answer Generation
w/o TE 0.6059 $\downarrow$ 17.98% 0.6482 $\downarrow$ 18.00%
w/o EA 0.6236 $\downarrow$ 15.58% 0.6555 $\downarrow$ 17.08%
w/o Proj 0.7212 $\downarrow$ 2.37% 0.7834 $\downarrow$ 0.90%
@@
4.6 Impact of Pre-trained SLM
One advantage of our framework is that it requires training only a small pre-trained language model (typically with fewer than 200 million parameters). We further investigate the impact of various pre-trained language models on overall performance. We selected four representative pre-trained models, including DistilBERT Sanh (2019) used in our EvidenceMap, and three larger models (BERT-Base Devlin et al. (2019), RoBERTa Liu (2019) and ModernBERT Warner et al. (2024)). We can observe from the Table 4 that:
@@ 4.6 Impact of Pre-trained SLM
w/ BERT-Base 0.7557 $\uparrow$ 0.7841 $\downarrow$
@@
w/ RoBERTa 0.7376 $\downarrow$ 0.7684 $\downarrow$
    - Base 125M
w/ ModernBERT 0.7425 $\uparrow$ 0.7970 $\uparrow$
@@
w/ BERT-Base 0.6672 $\downarrow$ 0.6526 $\downarrow$
@@
w/ RoBERTa 0.6614 $\downarrow$ 0.6547 $\downarrow$
    - Base 125M
@@
Table 4: Experiment on the impact of using different small pre-trained language models. DistilBERT, despite having the fewest parameters, demonstrates strong capabilities in learning evidence analysis for question answering.
Despite having the fewest parameters, DistilBERT remarkably achieves strong performance. DistilBERT retrains most of the ability of BERT but is much lighter, which aligns well with our framework, further illustrating the training of a suitable pre-trained SLM can produce better results than just selecting larger models in our framework.
The most powerful pre-trained language model can still achieve the best results. ModernBERT, the latest encoder-only model with a larger parameter size and higher efficiency compared to older models, achieves the best results on both datasets. This indicates that as SLM continues to update and advance, the capabilities of our framework will also be further enhanced.

### v2

# EvidenceMap: Learning Evidence Analysis to Unleash the Power of Small Language Models for Biomedical Question Answering

@@ Abstract
When addressing professional questions in the biomedical domain, humans typically acquire multiple pieces of information as evidence and engage in multifaceted evidence analysis to provide high-quality answers. Current LLM-based answer generation methods lack a detailed definition and learning process for evidence analysis, leading to the risk of error propagation and hallucinations while using evidence. Although increasing the parameter size of LLMs can alleviate these issues, it also presents challenges in model training and deployment with limited resources. In this study, we propose EvidenceMap, which aims to enable a tiny pre-trained language model to explicitly learn multiple aspects of biomedical evidence, including supportive evaluation, logical correlation and content summarization, thereby latently guiding a small generative model (around 3B parameters) to provide textual responses. Experimental results demonstrate that our method, fine-tuning a language model with 66M parameters, exceeds the RAG method with an 8B LLM by 19.9% and 5.7% in reference-based quality and accuracy, respectively.
EvidenceMap: Learning Evidence Analysis to Unleash the Power of Small Language Models for Biomedical Question Answering
Chang Zong1, Jian Wan1, Siliang Tang2, Lei Zhang1 thanks: Corresponding author. 1Zhejiang University of Science and Technology 2Zhejiang University zongchang@zust.edu.cn
@@ 1 Introduction
Answering specialized questions typically requires gathering a wide range of information as evidence, such as relevant texts from academic articles and domain knowledge Louis et al. (2024); Singhal et al. (2025). In the biomedical domain, it is crucial to formulate answers by evaluating and summarizing various evidence Singhal et al. (2023); Zhang et al. (2024). Figure 1 depicts a modern scenario in which a biomedical expert acquires relevant evidence from PubMed papers and a large language model (LLM) and performs a series of analytical steps that include supportive evaluation Jin et al. (2021), logical correlation Chen et al. (2021), and content summarization Xie et al. (2022), ultimately leading to a high-quality answer.
Figure 1: A diagram illustrates the evidence analysis process when addressing biomedical questions.
Generative question answering based on LLM has shown advantages in recent years. Current methods primarily leverage external knowledge for implicit reasoning Lewis et al. (2020); Asai et al. (2023); He et al. (2024); Wang et al. (2024) or utilize the ability of LLMs to generate explicit thinking processes Wei et al. (2022); Lightman et al. (2023); Yao et al. (2024), thus mitigating hallucinations and producing accurate and comprehensive responses. However, as illustrated in Figure 2, these approaches still encounter error propagation when confronted with various evidence due to lack of explicit analysis learning. Moreover, although increasing the scale of model parameters can improve reasoning capabilities, it poses challenges for training and deployment under limited resources.
Figure 2: Existing methods (left) suffer from unreasonable use or incorrect reasoning on low-quality evidence. Our framework (right) reduces error propagation by learning evidence analysis with small language models.
To alleviate the aforementioned issues, this study focuses on answering biomedical questions with various evidence and proposes a novel framework, named EvidenceMap. The framework explicitly learns and utilizes evidence analysis through small language models, thereby efficiently mimicking human problem-solving processes. Specifically, our framework models an evidence map using multiple pieces of evidence relevant to the question, and subsequently employs a small pre-trained language model (PLM) to learn the representation of components in evidence analysis including supportive evaluation, logical correlation, and content summarization, which serve as the analytical prompt to drive a small generative language model (GLM) to autoregressively derive answers. The experimental results demonstrate that applying evidence analysis to small language models significantly improves biomedical question-answering performance.
@@
  • We present EvidenceMap, a novel question answering framework inspired by biomedical evidence analysis, which implements multifaceted analysis including supportive evaluation, logical correlation, and content summarization in multiple pieces of evidence.
  • We develop an efficient evidence analysis learning method that utilizes semantic abilities of small pre-trained language models to obtain representations of various components of analysis, thereby guiding generative language models to produce the final answer.
  • Extensive experiments and analyses on biomedical datasets demonstrate that our approach, by explicitly learning and utilizing evidence analysis, achieves better performance on smaller models compared to larger-scale reasoning methods.
@@ 2.1 Evidence Map Modeling
We construct an evidence map ($\mathcal{M}$) for each input question to facilitate reasoning during generation. An evidence map introduces three key types of information, including the evaluation of how the evidence supports the question ($\mathcal{R}_{eval}$), the correlation between the evidence ($\mathcal{R}_{cor}$), and the summarization of all the evidence ($E_{sum}$). Generally, an evidence map can be formulated as follows.
@@
where $Q$ is the question to be solved, $Agg(\cdot)$ is the aggregation function to derive the summarization $E_{sum}$, $\mathcal{E}_{s}$ represents a set of evidence from the source $s$, which is denoted as follows.
@@
2.2 Learning Generative QA with LMs
@@
where $f_{plm}(\theta^{*})$ is a tiny pre-trained model (e.g. DistilBERT-Base), $f_{glm}$ is a small generative model (e.g. Llama-3.2-3B) for providing the best answer $A^{*}$. Then, our learning process is to find the optimized parameter set $\theta^{*}$, which leads to a minimum generative loss for all question-answer pairs, given the corresponding evidence map $EM_{i}$ for each question $Q_{i}$. This can be denoted as below.
@@ 3 Framework
Our EvidenceMap framework is illustrated in Figure 3, in which a pre-trained language model is fine-tuned to learn multifaceted evidence analysis to facilitate responses with a small generative model. The three main steps are described below.
@@ 3.1 Evidence Acquisition
As described in Figure 1, we gather relevant evidence from multiple sources for each question, including the internal knowledge of an LLM and the external information from academic papers. Specifically, for the evidence of an LLM, we ask GPT-4o Hurst et al. (2024) to take advantage of its knowledge to provide a series of key points that support the question, described in a concise language (see Appendix A.2). In addition, we extract paper snippets from existing biomedical QA datasets such as BioASQ Krithara et al. (2023) and PubMedQA Jin et al. (2019), which have already been annotated by human experts, as evidence related to a question. Therefore, an evidence-augmented QA dataset is constructed, in which each question corresponds to one piece of LLM evidence $E_{lm}$ and multiple pieces of paper evidence from $E_{p}^{1}$ to $E_{p}^{n}$. The input data can be described as follows.
@@
3.2 LM-based Evidence Analysis
Evidence Map.
As defined in Section 2.1, each input question, along with its associated evidence, can be modeled as an evidence map to facilitate reasoning. In addition to the evidence itself, an evidence map also incorporates multifaceted information on analysis, including the evaluation of how each evidence supports the question $R_{eval}^{i}$, the correlation between any two pieces of evidence $R_{cor}^{(i,j)}$, and the summarization of all the evidence $E_{sum}$.
Evidence Analysis.
We prompt a small pre-trained language model to perform the multifaceted evidence analysis in latent space. Specifically, for evidence summarization, we first obtain the overall semantics of each evidence node, and then use a multilayer perceptron (MLP) to aggregate and produce a summarized representation of the evidence, which is formulated as below.
@@ 3.2 LM-based Evidence Analysis
where $X$ is the concatenated feature of all evidence nodes, denoted as $[X_{lm},X_{p}^{1},\cdots,X_{p}^{n}]$, $H_{s}^{(n)}$ is the embedding of the $n$-th layer. Inspired by PromptBERT Jiang et al. (2022) and PromptEOL Jiang et al. (2024), we prompt the pre-trained language model $f_{plm}$ and get the last hidden states of the model as the evidence feature $X_{p}^{i}$, which is described as follows.
@@
Meanwhile, the evaluation of how an evidence node supports the question $Q$ is derived from the same language model by using $\texttt{Prompt}_{e}$ as below.
@@ 3.3 Analysis-Augmented Generation
With the result of the analysis by handling an evidence map, the final answer to the question is provided in an autoregressive way using a generative language model $f_{glm}$. We first convert all textual evidence and the question into an embedding $H_{t}$ using a text embedder, which is the first embedding layer of $f_{glm}$ as follows.
@@ 3.4 Training Objective
The autoregressive training objective concentrates on updating the parameter set $\theta$ of a small language model to provide the representation of evidence analysis to predict subsequent tokens accurately by a generative model. We calculate the probability of generating the $i$-th token of the target answer $A$ with the overall prompt and all previous tokens as follows.
@@ 4.1 Implementation Details
In the step of learning evidence analysis, we employ DistilBERT Sanh (2019), a lighter but faster model distilled from BERT Devlin et al. (2019), as the pre-trained model to be updated. In the generation step, the 1B and 3B versions of Llama-3.2 Dubey et al. (2024) are used to provide answers. We conduct experiments on two popular biomedical datasets, BioASQ Krithara et al. (2023) and PubMedQA Jin et al. (2019). During training, we set the learning rate as 5e-4, the batch size as 4, and the maximum number of paper evidence per question as 5. For each experiment, we train the language model for 10 epochs in a single NVIDIA A6000 GPU and evaluate the model on test sets.
@@ 4.3 Main Results
4o-mini-RAG - 0.1549 0.5848 0.7039 0.1328 0.5501 0.5874
Inference-LLM Llama3.1-RAG 8B 0.1936 0.6077 0.7101 0.1835 0.6431 0.6440
OpenBioLLM 8B 0.2196 0.5981 0.7355 0.1936 0.5726 0.6276
Llama3.2-RAG 3B 0.2956 0.6225 0.6843 0.1942 0.6102 0.6361
@@
Tuning-SLM RAG-LoRA 3B 0.3987 0.6834 0.7240 0.2443 0.6454 0.6342
EvidenceMap 1B 0.4245 0.7175 0.7311 0.2548 0.6513 0.6375
EvidenceMap 3B 0.4429 0.7287 0.7509 0.2931 0.6783 0.6610
@@
Second, Learning evidence analysis unleashes the power of small language models to efficiently resolve biomedical questions. Compared to larger models, including the RAG method with general LLMs (GPT-4o-mini, Llama-3.1-8B) and biomedical LLM (OpenBioLLM-8B), our approach demonstrates superior performance even with a 1B model. This indicates that evidence analysis can effectively unleash the capabilities of small language models.
@@ 4.4 Ablation Study on Evidence Analysis
w/o Eval 0.7085 $\downarrow$ 2.78% 0.7326 $\downarrow$ 2.44%
w/o Cor 0.6977 $\downarrow$ 4.25% 0.7270 $\downarrow$ 3.18%
w/o Sum 0.7142 $\downarrow$ 1.99% 0.7463 $\downarrow$ 0.61%
@@ 4.5 Ablation Study on Answer Generation
w/o TE 0.5977 $\downarrow$ 17.98% 0.6156 $\downarrow$ 18.02%
w/o EA 0.6152 $\downarrow$ 15.58% 0.6227 $\downarrow$ 17.07%
w/o Proj 0.7115 $\downarrow$ 2.36% 0.7442 $\downarrow$ 0.89%
@@
4.6 Impact of Pre-trained Language Models
One advantage of our framework is that it requires training only a small pre-trained language model to efficiently acquire evidence analysis abilities. We further investigate the impact of various pre-trained language models on overall performance. We selected four representative pre-trained models, with training strategies ranging from masked to casual style, including DistilBERT Sanh (2019) used in our EvidenceMap, and three larger models (BERT-Base Devlin et al. (2019), GPT2 Radford et al. (2019) and ModernBERT Warner et al. (2024)). We can observe from the Table 4 that:
@@ 4.6 Impact of Pre-trained Language Models
w/ BERT-Base 0.7457 $\uparrow$ 0.7441 $\downarrow$
@@
w/ GPT2 0.7299 $\uparrow$ 0.7576 $\uparrow$
    - Base 117M
w/ ModernBERT 0.7325 $\uparrow$ 0.7610 $\uparrow$
@@
w/ BERT-Base 0.6872 $\uparrow$ 0.6626 $\uparrow$
@@
w/ GPT2 0.6814 $\uparrow$ 0.6647 $\uparrow$
    - Base 117M
@@
Table 4: Experiment on the impact of using different small pre-trained language models. DistilBERT, despite having the fewest parameters, still demonstrates competitive capabilities in learning evidence analysis for question answering.
More powerful pre-trained language models can achieve better results. BERT-Base, GPT2, and ModernBERT, which have larger parameter sizes and better performance in general language abilities, improve results on both datasets. This indicates that as the language model continues to update and advance, the capabilities of our framework will also be further enhanced.
Despite having the fewest parameters, DistilBERT remarkably achieves competitive performance. DistilBERT retrains most of the ability of BERT but is much lighter, which fits well with our framework. With a reduction in parameter size, there is only a slight decline in performance, further illustrating that the evidence analysis we propose is easy to learn in small-scale models.